# and symmetric racks

### v1

@@ 2.5. Kink-involutory racks
We introduce kink-involutory racks, a class of racks that include quandles and involutory racks; cf. Figure 1.
@@
$s_{1}=(12)(345),\quad s_{2}=(12)(345),\quad s_{3}=(12),\quad s_{4}=(12),\quad\text{and }s_{5}=(12)$
@@ 4.2.1. Functoriality of $F$ and $F^{-1}$
For all $x\in X$, the fact that $\mathtt{u}$ and $s_{x}$ commute implies that
$\theta^{2}_{F(L)}(x)=\theta_{F(L)}\mathtt{u}s_{x}(x)=\theta_{F(L)}\mathtt{u}(x)=\mathtt{u}^{2}s_{x}(x)=x$
because $(X,s)$ is a quandle. Thus, $F(L)$ is kink-involutory, as desired. Moreover, if $L$ is a trivial Legendrian rack, then $F(L)=(X,s)$ is the underlying quandle of $L$, as desired.
@@ 5.2.2. The group $U_{R}$ of GL-structures on a rack
$U_{R}=C_{\operatorname{Aut}R}(\operatorname{Inn}R).$
@@ 5.3. Equivalence of symmetric kei and Legendrian kei
By comparing Definitions 3.3 and 5.4, we immediately deduce that $\mathsf{LK}$ and $\mathsf{SK}$ share exactly the same morphisms. Hence, $\mathsf{LK}$ and $\mathsf{SK}$ are isomorphic. ∎
@@
The category of involutory symmetric racks is isomorphic to the full subcategory of $\mathsf{LR}$ whose objects are Legendrian racks $(R,\mathtt{u})$ where $R$ is involutory and $\mathtt{u}$ is an involution.

### v2

@@ 2.5. Kink-involutory racks
We introduce kink-involutory racks, a class of racks that includes quandles and involutory racks; cf. Figure 1.
@@
$s_{1}=s_{2}=(12)(345),\quad\quad s_{3}=s_{4}=s_{5}=(12)$
@@ 4.2.1. Functoriality of $F$ and $F^{-1}$
For all $x\in X$, the fact that $\mathtt{u}$ is a GL-structure implies that
$\theta^{2}_{F(L)}(x)=\theta_{F(L)}\mathtt{u}s_{x}(x)=\theta_{F(L)}\mathtt{u}(x)=\mathtt{u}^{2}s_{x}(x)=x,$
where we have used the assumption that $(X,s)$ is a quandle. Thus, $F(L)$ is kink-involutory, as desired. Moreover, if $L$ is a trivial Legendrian quandle, then $F(L)=(X,s)$ is the underlying quandle of $L$, as desired.
@@ 5.2.2. The group $U_{R}$ of GL-structures on a rack
$U_{R}=C_{\operatorname{Aut}R}(\operatorname{Inn}R)\trianglelefteq\operatorname{Aut}R.$
@@ 5.3. Equivalence of symmetric kei and Legendrian kei
By comparing Definitions 3.3 and 5.4, we immediately deduce that the hom-sets in $\mathsf{LK}$ are exactly the same as the corresponding hom-sets in $\mathsf{SK}$. Hence, $\mathsf{LK}$ and $\mathsf{SK}$ are isomorphic. ∎
@@
The category of involutory symmetric racks is isomorphic to the full subcategory of $\mathsf{GLR}$ whose objects are GL-racks $(R,\mathtt{u})$ where $R$ is involutory and $\mathtt{u}$ is an involution.